OpenStudy (anonymous):

Find the axis of the graph of the function. f(x) = 4x2 + 8x + 7

OpenStudy (anonymous):

@jim_thompson5910 can u help plz!?

OpenStudy (anonymous):

y assume you mean x axis? in that case use quadratic formula.

OpenStudy (anonymous):

yeah i mean the x axis. does that mean just plug it into the x axis and that will b my answer? @timo86m

OpenStudy (anonymous):

it means you use the quadratic formula. Oh there is also an axis of symetry i think.

jimthompson5910 (jim_thompson5910):

I think you mean x-intercepts

OpenStudy (anonymous):

i meant to say plug it into the quadratic formula sorry!!!!

OpenStudy (anonymous):

x intercepts yes :)

jimthompson5910 (jim_thompson5910):

So that would mean you do this \[\Large x = \frac{-b\pm\sqrt{b^2-4ac}}{2a}\] \[\Large x = \frac{-(8)\pm\sqrt{(8)^2-4(4)(7)}}{2(4)}\] I'll let you finish up

OpenStudy (anonymous):

thanks guys!!!